OpenStudy (anonymous):

1) Identify the correct slope and y-intercept of the equation y - 4x = -3 A)slope = -1; y-intercept = (0, -3) B) slope = -1 over 4 ; y-intercept = (0, 3) C) slope = 4; y-intercept = (0, -3) D) slope = 4; y-intercept = (0, 3) 2) Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. A) y = 1 over 3x - 8 B) y = -3x + 12 C) y = 3x - 24 D) y = -1 over 3x - 4 just need help with these two then I can do the rest on my own.

jimthompson5910 (jim_thompson5910):

# 1 Solve for y: y - 4x = -3 y - 4x+4x = -3+4x y = 4x-3 The equation is now in slope intercept form y = mx+b where m is the slope and (0, b) is the y-intercept We can see that m = 4 and b = -3 So the slope is 4 and the y-intercept is (0, -3) So the answer is choice C ======================================================================== ======================================================================== # 2 Anything perpendicular to y = 3x - 6 will have a slope of -1/3. Simply flip the original slope 3/1 to get 1/3 and change the sign to get -1/3 So use this along with the point (6,-6) with y = mx+b to find the perpendicular equation You'll be plugging in m = -1/3 x = 6 y = -6 \[\Large y = mx+b\] \[\Large -6 = \left(\frac{-1}{3}\right)\left(6\right)+b\] \[\Large -6 = \frac{-6}{3}+b\] \[\Large -6 = -2+b\] \[\Large -6+2 = b\] \[\Large -4 = b\] \[\Large b = -4\] So the equation of the perpendicular line is \[\Large y = -\frac{1}{3}x-4\] So the answer is choice D

OpenStudy (anonymous):

Thank you, that helped me a lot, made more sense!

jimthompson5910 (jim_thompson5910):

Glad to be of help.

jimthompson5910 (jim_thompson5910):

and I'm glad things are starting to click for you.